four more than 2/5 of a number is 24.
Answer by stanbon(75887)

four more than 2/5 of a number is 24.
--------
(2/5)x + 4 = 24
--------
(2/5)x = 20
x = 100/2
x = 50
------------
Cheers,
Stan H.
----------